OpenStudy (anonymous):

a coin is tossed 5 times. whats the probability of it landing all on tails in a row?

OpenStudy (campbell_st):

P(T) = 1/2 5 in a row = (1/2)^5 = 1/32

OpenStudy (anonymous):

okkk thanks i was braindead today

OpenStudy (anonymous):

5c5 (1/2)^5 (1/2)^0